# Toward a Deterministic Nucleation Theory for Chirality-Controlled Nanotube Synthesis

### v2

@@ Conclusion
Through the discovery of the vector sum rule and its application to six-fold symmetric nucleation caps, we have reframed chirality-controlled nanotube synthesis as a deterministic process governed by cap topology and epitaxial matching with catalyst surfaces. While previous theories attributed chirality-selective growth primarily to edge matching mechanisms [36], our findings establish that nucleation is deterministically controlled by these topological and interfacial factors. This theoretical framework coherently explains the experimental enrichment of $(12,6)$ nanotubes while accounting for the emergence of chiral enrichment in neighboring species. The parallel enrichment observed around $(6,6)$ nanotubes—which similarly feature six-fold symmetric caps—further validates this mechanism [37]. Although our work captures the essential features of nucleation, developing predictive synthesis requires a more comprehensive understanding of how catalyst morphology directs pentagon formation. Our results provide the theoretical foundation for advancing toward this goal.
Acknowledgements
Fundings: This study was supported by the National Natural Science Foundation of China (Nos. 12462009 and 12132008). We thank Nader Anani for advice on the manuscript.
Data avalibility: All data are available in the main text or the supplementary materials. A direct implementation of the topological framework into computer code for designing carbon caps is available at GitHub (https://github.com/ZhengrongGuoChina/Graphene-topology).

### v3

@@ Conclusion
Through the discovery of the vector sum rule and its application to six-fold symmetric nucleation caps, we have reframed chirality-controlled nanotube synthesis as a deterministic process governed by cap topology and epitaxial matching with catalyst surfaces. While previous theories attributed chirality-selective growth primarily to edge matching mechanisms [35], our findings establish that nucleation is deterministically controlled by topological and interfacial factors. This framework coherently explains the experimental enrichment of $(12,6)$ nanotubes while accounting for chiral enrichment in neighboring species. The parallel enrichment observed around $(6,6)$ nanotubes—which similarly feature six-fold symmetric caps—further validates this mechanism [36]. Although our work captures essential nucleation features, developing predictive synthesis requires a more comprehensive understanding of how catalyst morphology directs pentagon formation. Our results provide the theoretical foundation for advancing toward this goal.
Acknowledgement
This study was supported by the National Natural Science Foundation of China (Nos. 12462009 and 12132008). We thank Nader Anani for advice on the manuscript.
Data, Materials, and Software Availability
All data are available in the main text or the supplementary materials. A direct implementation of the topological framework into computer code for designing carbon caps is available at GitHub (https://github.com/ZhengrongGuoChina/Graphene-topology).
Method
The formation energies in Fig. 3D were calculated using the LAMMPS package. Carbon-carbon interactions were described by the reactive bond-order (REBO) potential, while carbon-copper substrate interactions were modeled using a Lennard-Jones potential ($\sigma=2.89$ Å, $\epsilon=16.2$ meV). All simulations were performed in the canonical (NVT) ensemble at 1 K with a rigid substrate